OpenStudy (anonymous):

A manufacturer of gas grills has daily production costs of C = 400 – 5x + 0.125x2, where x is the number of gas grills produced. How many units should be produced each day to yield a minimum cost?

OpenStudy (amoodarya):

 hint : it is a quadratic function , so find minimum by finding c'(x) c'(x)=0 x_min is the root(s) of c'=0

OpenStudy (anonymous):

\[C = 400 – 5x + 0.125x^2,\] right?

OpenStudy (anonymous):

minimum is the second coordinate of the vertex first coordinate of the vertex is always \[-\frac{b}{2a}\] which in your case is \[-\frac{-5}{2\times .125}=20\]

OpenStudy (anonymous):

oh they just want the number to make, that answer is 20